OpenStudy (neer2890):

It's an aptitude question: Take five 3's and use any mathematical operation. your answer should be 31?

OpenStudy (shadowlegendx):

Are you limited to using only one type of mathematical operation?

OpenStudy (neer2890):

no, you can use any mathematical operation between two 3's

OpenStudy (shadowlegendx):

\[3 + 3 + 3 + 3 + 3 = 15 \] \[3 - 3 - 3 - 3 - 3 - 3 = -15 \] \[3 \times 3 \times 3 \times 3 \times3 = 243 \] \[3\div3\div3\div3\div3\div3 = 0.01234567901 \] Do these follow the rules?

OpenStudy (neer2890):

no but instead you can use 3+3-3*3/3 any mathematical operation between any two 3's

OpenStudy (shadowlegendx):

Mhm ok

OpenStudy (dan815):

3^5/x=31

OpenStudy (shadowlegendx):

@dan815 use 3's

OpenStudy (dan815):

5/log 3 base 31 =x

OpenStudy (neer2890):

you have to use only five 3's,not less than that and not more than that

OpenStudy (dan815):

(3*3*3*3*3)^(5/5/log 3 base 31)

OpenStudy (dan815):

:p

OpenStudy (dan815):

thats an operation on 5 3s

OpenStudy (neer2890):

but you used 5 and base 31

OpenStudy (dan815):

why cant that be an operation!!

OpenStudy (dan815):

fine ill do it properly!

OpenStudy (dan815):

how about 3^0 is that allowed

OpenStudy (neer2890):

no 3^3 is allowed

OpenStudy (neer2890):

operand should only be 3,operations can be any mathematical operations

hartnn (hartnn):

\(\huge 3^3 + \dfrac{3}{3} +3 \)

OpenStudy (dan815):

3^3+3+3/3

OpenStudy (dan815):

NO :(

hartnn (hartnn):

thought process: 3^3 =27 how do i get 3 using 3 threes how do i get 1 using 2 3s simple

OpenStudy (dan815):

foiled by hatnn

OpenStudy (dan815):

yes you had to have known u need a 1 no matter what because, 31 is not a factor of 3

OpenStudy (neer2890):

thanks @hartnn and @dan815

OpenStudy (shadowlegendx):

Hartnn too gewd c;

hartnn (hartnn):

***how do i get 4 using 3 threes

hartnn (hartnn):

guys, try 40 using 5 threes.

hartnn (hartnn):

@ShadowLegendX @dan815

OpenStudy (neer2890):

3+3/3=4 @hartnn

hartnn (hartnn):

oh i asked 40 not 4 that too using 5 threes, not more, not less

OpenStudy (anonymous):